OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that goes through the point 5,-6 with m=-3/2

OpenStudy (shamil98):

y = mx + b plug in your values , solve for b. put b and your slope back into the now finished equation to get slope-intercept form.

OpenStudy (shamil98):

(5,-6) --> (x,y) m = -3/2 y = mx +b -6 = (-3/2)(5) + b and solve

OpenStudy (anonymous):

3/2?